OpenStudy (anonymous):

Write an equation of the sine function with amplitude 3, period 3pi/2, and phase shift pi/4.

OpenStudy (anonymous):

So far I have y = sin (4/3x-c) Is this right and what do I do next?

OpenStudy (anonymous):

@jim_thompson5910 @kropot72 hey think one of you can help me?

jimthompson5910 (jim_thompson5910):

you have the correct equation that has a period of 3pi/2 now you need to deal with the amplitude and phase shift

jimthompson5910 (jim_thompson5910):

y = A*sin(Bx - C) + D |A| is the amplitude T = 2pi/B is the period C/B is the phase shift y = D is the midline

OpenStudy (anonymous):

Oops. I thought I put the amp in, I guess I didn't click the 3 hard enough lol. y = 3 sin (4/3x-c)

jimthompson5910 (jim_thompson5910):

I gotcha, so now you have the equation with the correct period and amplitude

jimthompson5910 (jim_thompson5910):

now you just need to deal with the phase shift

OpenStudy (anonymous):

(pi/4)/(4/3)?

jimthompson5910 (jim_thompson5910):

y = A*sin(Bx - C) + D |A| is the amplitude T = 2pi/B is the period C/B is the phase shift y = D is the midline

jimthompson5910 (jim_thompson5910):

C/B is the phase shift C is unknown right now B = 4/3

jimthompson5910 (jim_thompson5910):

C/B = pi/4

OpenStudy (anonymous):

C/(3pi/2)=pi/4?

jimthompson5910 (jim_thompson5910):

B is 4/3, not 3pi/2

OpenStudy (anonymous):

Duh...

OpenStudy (anonymous):

C/4/3=pi/4

jimthompson5910 (jim_thompson5910):

C/(4/3) = pi/4, correct solve for C

OpenStudy (anonymous):

C= pi/3?

jimthompson5910 (jim_thompson5910):

A = 3 B = 4/3 C = pi/3 D = 0

jimthompson5910 (jim_thompson5910):

\[\Large y = A\sin\left(Bx - C\right)+D\] \[\Large y = 3\sin\left(\frac{4}{3}x - \frac{\pi}{3}\right)+0\] \[\Large y = 3\sin\left(\frac{4}{3}x - \frac{\pi}{3}\right)\]

jimthompson5910 (jim_thompson5910):

D = 0 because by default, the midline is the x axis (when y = 0)

OpenStudy (anonymous):

Thank you so much. I get it now. That was confusing.

jimthompson5910 (jim_thompson5910):

yeah it's a bit weird and tricky, but you'll get the hang of it all